OpenStudy (anonymous):

:O PLEASE HELP The current, I , in amperes, that flows through a coil of wire is found by dividing 110 by the resistance of the wire, R, in ohms. Using I(R)=100/R, find the current through a coil of wire with a resistance of 80 ohms.

OpenStudy (anonymous):

I(R)=100/R or I(R)=110/R ?

OpenStudy (anonymous):

I(R)=110/R, I'm sorry :/

OpenStudy (anonymous):

plug in 80 for r I(80)=110/80=11/8= 1.375 amps

OpenStudy (anonymous):

O.k, thanks! Do you know what that number would be in fraction form?

OpenStudy (anonymous):

i wrote it in fraction form and decimal form above

OpenStudy (anonymous):

I meant 1.375

OpenStudy (anonymous):

11/8=1.375

OpenStudy (anonymous):

Oh wait. I see it. :/ THANKS